OpenStudy (baldymcgee6):

Describe what the level curve f(x,y) =2 for 2 = \frac{\sqrt{x-y}}{\sqrt{x+y}} looks like graphically

OpenStudy (baldymcgee6):

\[2 = \frac{\sqrt{x-y}}{\sqrt{x+y}}\]

OpenStudy (anonymous):

\[4=\frac{ x-y }{ x+y } \]level curve ?

OpenStudy (baldymcgee6):

I think the one I posted is already the level curve.

OpenStudy (baldymcgee6):

\[f(x,y) = \frac{{\sqrt {x - y} }}{{\sqrt {x + y} }}\] This is the original function

OpenStudy (anonymous):

according to your equation\[4x+4y=x-y\] \[3x+5y+2=2\] \[f(x,y)=3x+5y+2\]

OpenStudy (anonymous):

since f(x,y)=2

OpenStudy (baldymcgee6):

This is the level curve \[2 = \frac{\sqrt{x-y}}{\sqrt{x+y}}\] isnt it?